OpenStudy (anonymous):

A board 12 3/4 feet long is cut so that 9 7/8 feet are left. What length of the board was cut off? A) 2 1/2 feet B) 2 7/8 feet C) 3 1/2 feet D)3 7/8 feet

OpenStudy (anonymous):

\[The~length~of~the~board~was~cut~off=12\frac{ 3 }{ 4 }-9\frac{ 7 }{ 8 }\]

OpenStudy (anonymous):

wait, do i just subtract?

OpenStudy (mrnood):

yes - just subtract

OpenStudy (anonymous):

ill tell u my answer, k?

OpenStudy (anonymous):

\[=\color{red}{12\frac{ 3~times~2 }{ 4~times~2 }}-9\frac{ 7 }{ 8 }\]make the denominator same by multiplying 2

OpenStudy (anonymous):

B

OpenStudy (anonymous):

im telling him what i got from HIS explanation..

OpenStudy (anonymous):

CLOSING QUESTION.

OpenStudy (mrnood):

oops sorry didint realise you were the poster!

OpenStudy (anonymous):

oh LOL its ok :)

OpenStudy (anonymous):

;)